Grullon Co. is considering a 7-for-3 stock split. The current stock price is $75.00 per share, and the firm believes that its total market value would increase by 5% as a result of the improved liquidity that it thinks would follow the split. What is the stock's expected price following the split?
Answer #1
The answer is E(x) = $33.75

Given: 5% increase
3shares x $75 = $225 * 1.05 = 236.25 / 7 shares = $33.75